Circular No. 7972

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2002 S1 B. Skiff, Lowell Observatory, reports his discovery of a comet on images

taken by him in the course of the LONEOS program, with the available observations given below. The object has a moderately condensed coma of diameter 10′′ and a tail 30′′ long in p.a. 285^{o}. At Skiff’s request, M. Hayes-Gehrke obtained a confirming 5-min CCD R frame (also on Sept. 17, in twilight) with the Lowell Observatory’s Perkins 1.8-m telescope, showing a 8′′ coma and a tail ∼ 30′′ long in p.a. 280^{o} or 285^{o}.

2002 UT  α_{2000}  δ_{2000}  m_{1}
Sept.17.41755  7^{h}02^{m}13.^{s}22  +49^{o}41′30′′.4  18.0
17.43083  7 02 14.77  +49 41 35.0
17.44419  7 02 16.21  +49 41 38.3
17.45753  7 02 17.87  +49 41 43.0

SUPERNOVA 2002fi IN MCG –04-7-10 D. Weisz and W. D. Li, University of California at Berkeley, report the

LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered KAIT CCD images taken on Sept. 13.4 (mag ∼ 18.5) and 17.4 UT (mag ∼ 18.3). SN 2002fi is located at α = 2^{h}33^{m}54^{s}.40, δ = −21^{o}02′29′′.4 (equinox 2000.0), which is 19′′.4 west and 6′′.7 north of the nucleus of MCG –04-7-10. A KAIT image taken on Aug. 28.4 showed nothing at this position (limiting mag ∼ 19.0).

V4742 SAGITTARII = POSSIBLE NOVA IN SAGITTARIUS N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V4742 Sgr has been given to this object (cf. IAUC 7971).

H. Yamaoka, Kyushu University, reports the following position for this object from a CCD frame taken by B. Monard on Sept. 16.7 UT: α = 18^{h}02^{m}21^{s}.864, δ = −25^{o}20′32′′.22 (equinox 2000.0; 102 UCAC1 reference stars); no source is visible on Digitized Sky Survey images within 3′′ of this position. C. Buil, Ramonville St. Agne, France, reports that a spectrum obtained on Sept. 16.806 yields an expansion velocity of 1100 ± 150 km/s from the FWHM of the Hα line, which shows a P-Cyg profile. T. Kato, Kyoto University, remarks that Buil’s spectrum appears to be that of a nova just prior to maximum. Visual magnitude estimates: Sept. 16.854, 8.0 (P. Schmeer, Bischmisheim, Germany); 17.040, 8.1 (R. Y. Shida, São Paulo, Brazil); 17.496, 8.7 (A. Pearce, Nedlands, W. Australia).

2002 September 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green